OpenStudy (anonymous):

The value of cosec (-750degee) is

OpenStudy (akshay_budhkar):

-2

OpenStudy (anonymous):

degree

OpenStudy (akshay_budhkar):

it is same as cosec 30

OpenStudy (anonymous):

how you got that ? can u explain?

OpenStudy (anonymous):

cosec (-750) =-cosec750 =-(cosec(360+390)) =-cosec(390) =-cosec(360+30)=-cosec30

OpenStudy (anonymous):

which is -2

OpenStudy (anonymous):

i hope u have understood

OpenStudy (anonymous):

no i couldn't

OpenStudy (akshay_budhkar):

cosec(360+x) is cosec x you know this?

OpenStudy (akshay_budhkar):

and cosec(-x)=-cosecx

OpenStudy (akshay_budhkar):

you know these formulae right?

OpenStudy (anonymous):

yes

OpenStudy (akshay_budhkar):

ya so thats what seikumar has used in his steps.. you couldnt get it?

OpenStudy (anonymous):

in 4th step?

OpenStudy (akshay_budhkar):

cosec (-750) =-cosec750 (formula 2) =-(cosec(360+390)) =-cosec(390) (formula 1) =-cosec(360+30) =-cosec30 (formula 1)

OpenStudy (akshay_budhkar):

you get it?

OpenStudy (anonymous):

ya

OpenStudy (akshay_budhkar):

kk great! keep it up!

OpenStudy (anonymous):

thanks

OpenStudy (akshay_budhkar):

ur welcome :)